OpenStudy (anonymous):

What is the slope line that passes through points (0,-2)(1 (1,4)

geerky42 (geerky42):

\[m = \dfrac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

That means y=6

geerky42 (geerky42):

yes. what does x = ?

geerky42 (geerky42):

we want m, not y.

OpenStudy (anonymous):

Oh 6/1

geerky42 (geerky42):

Yes, so m = 6/1 = \(\boxed{6}\) Is that clear?

OpenStudy (anonymous):

Ya